# MoveTouch: Robotic Motion Capturing System with Wearable Tactile Display to Achieve Safe HRI

### v1

@@ Abstract
The collaborative robot market is flourishing as there is a trend towards simplification, modularity, and increased flexibility on the production line. But when humans and robots are collaborating in a shared environment, the safety of humans should be a priority. We introduce a novel wearable robotic system to enhance safety during Human-Robot Interaction (HRI). The proposed wearable robot is designed to hold a fiducial marker and maintain its visibility to the tracking system, which, in turn, localizes the user’s hand with good accuracy and low latency and provides haptic feedback on the user’s wrist. The haptic feedback guides the user’s hand movement during collaborative tasks in order to increase safety and enhance collaboration efficiency. A user study was conducted to assess the recognition and discriminability of ten designed haptic patterns applied to the volar and dorsal parts of the user’s wrist. As a result, four patterns with a high recognition rate were chosen to be incorporated into our system. A second experiment was carried out to evaluate the system integration into real-world collaborative tasks.
@@
Haptic feedback Human Robot Interaction Wearable devices.
@@ 1 Introduction
The number of robots in the industry is increasing globally, reaching over 3.9 million robots in factories in 2022, according to the International Federation of Robotics (IFR) [8]. Robots are considered helpful associates for carrying out repetitive tasks and operating in hazardous environments, while human dexterity can be harnessed in the operation. However, robots are mostly being implemented in the industry as tools, not as companions for humans. To achieve collaboration between humans and robots, the safety of users has to be guaranteed [21].
Haptic feedback is one of the most efficient ways to guarantee safety [18], which is applied in several domains, including medical applications [6], virtual reality [4], and assisting blind people in indoor environments [9]. All these previous studies are concerned with providing haptic feedback to ensure the safety of users during a hazardous scenario and encouraging people to perform an act to avoid dangerous situations.
Tactile feedback improves mutual perception during human-robot collaboration by communicating the robot’s motion intentions and its status to a human worker using hand-worn haptic feedback devices [7]. Several research endeavors have focused on enhancing safety in HRI scenarios through haptic devices [13]. Haptic feedback works as an excellent notification system for humans [20]. Scheggi et al. attempted to use haptic feedback for human navigation with a mobile robot through a vibrotactile bracelet [17].
Tactile feedback can be given to users on different parts of the body, for example, on the forearm [11], [1], [3], or the palm [5]. Evarsson et al. investigated the perceptual vibrotactile thresholds for a range of frequencies on both the inside and outside areas of the wrist, showing that personalized threshold measurements at the actuator locations will be required in order to fine-tune a device for the user [1]. Stanley et al. evaluated ten forms of tactile feedback using five wearable actuators, showing that repeated taps on the subject’s wrist on the side toward which they should turn enhanced the best performance of the subjects [19].
The main focus of this work is to study different haptic patterns on the up (dorsal) and down (volar) parts of the wrist. We also introduce an improved version of our robotic device, dubbed ArUcoGlide that was previously published [2]. We have enhanced the design of the ArUcoGlide making the device smaller in size based on a differential gear train mechanism.
@@ 2 MoveTouch Motion Capture System
    A wearable robot, dubbed MoveTouch, is responsible for both adjusting the orientation of an ArUco marker held at its end-effector and providing haptic feedback to the user. The movable ArUco marker ensures the visibility of the marker to the motion capture system, avoiding any occlusion due to the movement of the user’s hand. The haptic feedback is provided to the user through five vibration motors controlled independently, which are able to generate different patterns. The haptic system acts as a guidance system that tells the user about the optimal movement of his hand during Human’Robot Interaction (HRI) to avoid any collisions.
@@ 2.1 MoveTouch Wearable Robot
The primary goal of the MoveTouch wearable robot is to assist in determining the user’s position and provide vibrotactile guidance, ensuring a secure (HRI) experience.
@@ 2.1.1 Mechanical Design
The proposed device is a two-degrees-of-freedom (2-DoFs) wearable robot, based on a differential gear train mechanism, that holds an ArUco marker at its end-effector with the ability to rotate this marker around two perpendicular axes through two servo motors [10], [12]. Additionally, it includes a bracelet equipped with vibration motors to deliver haptic patterns to the user. The links, holders, and gears were designed and 3D-printed with PLA material; the 3D CAD model is shown in Fig. 1. The device maintains the visibility of the ArUco marker to the tracking system by continuously adjusting the motors’ angular position to hold the marker in a fixed orientation to the camera frame.
Figure 1: 3D CAD model of the MoveTouch wearable device.
With the aim of generating two independent motions by means of identical motors placed beside each other, we decided to use a differential gear train mechanism as shown in Fig. 1(a). In this figure, the fixed base, the differential gear train, and the rotating Aruco holder are shown.
We denote $\Delta\theta_{A}$ and $\Delta\theta_{B}$ as the rotation angles of the first and second motors, respectively. The rotation angles of the ArUco marker around the two axes are $\Delta\theta_{P}$ and $\Delta\theta_{C}$ which can be seen in Fig. 1(b). The rotations from motors A and B are transmitted to the left and right sun gears of the differential unit. The rotation angles are related by equations:
$\Delta\theta_{A}=n_{a}(n_{s}\Delta\theta_{P}+\Delta\theta_{C}),$ (1)
@@
where $n_{a}=n_{b}=0.5$ are the gear ratio of the motor gears to the sun gears and $n_{s}=0.5$ is the gear ratio of the sun gears to the planet gear.
We chose to integrate five vibration motors within the system. The selection of this number takes into consideration two key factors: 1) Enabling a uniform distribution of the vibration motors on one side of the user’s wrist (either dorsal or volar), which typically falls within the range of 15-20 cm. 2) This number of vibration motors is sufficient for generating a diverse array of vibrotactile patterns to be employed in user experiments.
@@ 2.1.2 Electronic Design
The electronic parts consist of an ESP32 microcontroller and two Gotech GS-9025MG servo motors that need a supply voltage of 4.8–6 V and give a minimum torque of 2.5 kg.cm. The system is powered by a Li-Po 7.4 V battery connected through a DC/DC converter to power the microcontroller and the servo driver.
Five coin vibration motors, able to vibrate at a frequency in the range 10 to 55 Hz, were located within a distance of 2 cm between them on the bracelet.
The position of the servo motors and the frequency of the vibration motors are controlled via Bluetooth from a base computer.
@@ 2.2 Motion capture system
To ensure the safety of humans, we need to track the real-time position of the operator within the working space. Our proposed tracking system is both cost-efficient and easy to install, utilizing an ArUco marker to track the operator’s hand. The system comprises a base computer and an HD webcam C930e from Logitech mounted on a stand that can be adjusted to capture different angles of the workspace, providing greater flexibility for the user. The basic idea of the tracking system is to transfer the user’s hand position into the UR10 robot base coordinate system so that we can avoid any collision with it. Since the user’s hand position is known in the camera coordinate system, the transformation from the marker to the camera $T_{C}^{A}$ is known. By attaching another Aruco marker (we will call it the base marker) at a known transformation from the UR10 robot base, we can also get the position of the base of the robot in the camera coordinate system; thus, we can get the transformation $T_{C}^{B}$. As a result, we obtain the transformation between the user’s hand and the UR10 robot base coordinate system $T_{B}^{A}$, and thus the position of the user’s hand in the UR10 robot base coordinate system as follows:
@@
Calculating the transformation from the base marker to the camera $T_{C}^{B}$ is required only once before starting the experiments, as the camera will be in the same position throughout the whole experiment. However, if the camera’s position or orientation is altered, the process needs to be repeated to derive the correct transformation matrix. Once the transformation matrix $T_{B}^{C}$ is determined, we can utilize it to track the position of the ArUco marker that is attached to the user’s hand, enabling us to locate the marker within the UR10 robot’s coordinate system.
@@ 3 Vibrotactile Guidance System
The proposed wearable device, MoveTouch, provides vibrotactile guidance, facilitating human-robot collaboration. This innovative system harnesses tactile sensations to guide users in adjusting their hand position when the collaborative robot approaches and alerts them about potentially risky situations. To ensure effective information transfer to the user, it’s crucial to understand how humans perceive tactile pattern sensations. This perception is achieved through mechanoreceptors, including Merkel cells, Meissner’s corpuscles, and Pacinian corpuscles, situated in the outer two layers of the skin—the epidermis and dermis. Merkel cells specialize in sensing sustained pressure and static touch. Meissner’s corpuscles are sensitive to low-frequency vibrations within the range of approximately 10 to 50 Hz [14], while Pacinian corpuscles are sensitive to higher-frequency vibrations, typically around 20 to 1000 Hz [15], [16].
@@ 3.1 Patterns Design
Given that our wearable robot is worn on the wrist of the user, it is prudent to incorporate vibrotactile patterns in this specific region. We need to first assess the user’s ability to recognize vibrotactile patterns on their wrist and determine suitable frequencies for these patterns.
Five different patterns were designed using the five vibration motors, each with two different frequencies. The tactile patterns are illustrated in Fig. 2.
@@ 3.2 Pattern Recognition
In this experiment, we aimed to assess the ability of the participant to recognize and differentiate between the designed vibrotactile patterns. We invited 12 participants and asked them to tell us which haptic pattern they perceived. Their data was recorded and analyzed.
@@ 3.2.2 User Study Procedure
During the study, each user was asked to wear the MoveTouch robot and sit in front of a PC that showed a graphical user interface (GUI) that allowed them to choose the pattern that they felt. An example of the user study of this experiment is shown in Fig. 3.
Figure 3: The experimental setup, a) the pattern recognition experiment setup, b) the system evaluation in a real collaborative task setup.
To identify the optimal wrist location for wearing the device, participants were evaluated twice, wearing the device each time on the upper or lower part of their wrists. The participants were split into two groups: the first group tried the device on the upper part first, followed by the lower part, while the second group followed the opposite sequence. Each of the 10 patterns was replicated five times blindly in a random order; thus, 50 patterns were provided to each participant in each evaluation. Participants were asked to wear headphones while playing white noise and engaged in the experiment through the GUI interface. They had the option to initiate the vibration pattern when they were ready and select from the displayed patterns the one that they felt. During the experiment, participants were not allowed to repeat any pattern.
@@ 3.2.4 Conclusion
The average recognition rate of all patterns on the dorsal part of the wrist was 66.6%, while on the volar side of the wrist, it was 75.64%. We can notice that almost all the patterns are better recognized on the lower part of the wrist; thus, this position of the vibration motors was selected. The patterns that presented a recognition rate higher than 80% were chosen: patterns 1L, 2L, 3L, and 5H will be used for further studies.
3.3 System evaluation in real collaborative task
In this evaluation, the system was integrated into a real collaborative task to study its effectiveness in enhancing the overall safety of the user and guiding him to a desired position. The four chosen patterns were integrated to guide the user on how to move his hand as follows: Pattern 1L means to move the hand to the right, pattern 2L means to move the hand to the left, pattern 3L means to move the hand down, and pattern 5H means to move the hand back.
@@ 3.3.2 Experimental setup
The participants were asked to stand in front of a Siegmund welding table, where a UR10 robot was performing a pick-and-place task of cubes positioned at specific locations on the table. The participants wore the MoveTouch on their right wrist and utilized an eye cover to eliminate any visual feedback about the location of the robot’s Tool Center Point (TCP). Participants were also performing a pick-and-place task in the same box that the robot was using. The surveillance camera was placed on a tripod near the table, where it could monitor the entire workspace. An observer was present during the experiment, poised to activate the emergency button to halt the robot in case of any unforeseen issues. The experimental setup is shown in Fig. 3.

### v2

@@ Abstract
The collaborative robot market is flourishing as there is a trend towards simplification, modularity, and increased flexibility on the production line. But when humans and robots are collaborating in a shared environment, the safety of humans should be a priority. We introduce a novel wearable robotic system to enhance safety during Human-Robot Interaction (HRI). The proposed wearable robot is designed to hold a fiducial marker and maintain its visibility to a motion capture system, which, in turn, localizes the user’s hand with good accuracy and low latency and provides vibrotactile feedback to the user’s wrist. The vibrotactile feedback guides the user’s hand movement during collaborative tasks in order to increase safety and enhance collaboration efficiency. A user study was conducted to assess the recognition and discriminability of ten designed vibration patterns applied to the upper (dorsal) and the down (volar) parts of the user’s wrist. The results show that the pattern recognition rate on the volar side was higher, with an average of 75.64% among all users. Four patterns with a high recognition rate were chosen to be incorporated into our system. A second experiment was carried out to evaluate users’ response to the chosen patterns in real-world collaborative tasks. Results show that all participants responded to the patterns correctly, and the average response time for the patterns was between 0.24 and 2.41 seconds.
@@
Haptic feedback Human Robot Interaction Wearable devices Motion Capture.
Figure 1: MoveTouch, a novel wearable robot for position tracking and haptic feedback. An ArUco marker is located at the end effector to adjust its orientation, while five vibration motors deliver tactile feedback to the user.
@@ 1 Introduction
The number of robots in the industry is increasing globally, reaching over 3.9 million robots in factories in 2022, according to the International Federation of Robotics (IFR) [10]. Robots are considered helpful associates for carrying out repetitive tasks, while human dexterity can be harnessed in the operation. However, robots are mostly being implemented in the industry as tools, not as companions for humans. To achieve collaboration between humans and robots, the safety of users has to be guaranteed [19].
Haptic feedback is considered one of the most efficient ways to guarantee safety in the application of Human-Robot Interaction (HRI) [16]. It works as an excellent notification system for humans [18].
Several studies have focused on enhancing safety in HRI scenarios through tactile devices [15]. Some have included tactile feedback in medical applications [7], virtual reality [4] or even in assisting blind people in indoor environments [11]. Tactile feedback can be applied to different parts of the human body, for example, to the fingertips [7], to the palm [5] or to the forearm [13], [3], [1].
In the applications where users need their hands to perform some tasks, providing tactile feedback to the forearm is considered a preferable option since it allows users to use their hands freely. Stanley et al. evaluated ten forms of tactile feedback to the human wrist using five wearable actuators (taper, dragger, squeezer, twister, and vibration) [17]. They showed that repeated taps on the subject’s wrist on the side toward which they should turn enhanced the performance of the subjects. However, their work compared between different types of tactile feedback based on the suggested tactile actuators and did not evaluate different patterns of vibration. Chase et al. showed that, by training, it is possible to improve the signal identification in haptic feedback for novices [6].
This research work showed that tactile feedback can enhance the human response and help users perform some actions depending on the haptic feedback that they receive. However, they did not evaluate the different patterns in terms of how effectively users can recognize these patterns. Hong J. et al. developed a wrist haptic device to guide blind people’s hands to perform different tasks [9]. They found, by using 4 and 8 different vibro motors, that single motor feedback was more efficient than interpolated feedback (using more than one motor at a time). However, they did not show the difference between giving feedback to the volar and the dorsal part of the forearm.
ArUcoGlide is a wearable robotic device that we previously developed to ensure that the motion capture system is always able to track the hand of the user [2]. The main idea of this device is to have a rotatable ArUco marker with two Degrees of Freedom (DOF) to maintain a specific angle with the camera and thus avoid any possible occlusion. In this work, we have enhanced the design of the ArUcoGlide to make it less bulky based on a differential gear train mechanism while ensuring the same performance and adding vibro motors to give tactile feedback to the user’s wrist.
This paper aims to study the perception of humans for different tactile patterns exerted on the up (dorsal) and down (volar) parts of the wrist. The tactile feedback is applied through five vibro motors attached to a wearable band. A human study was conducted to compare between the applied patterns depending on the position (dorsal or volar) and the applied frequency. Following the results of the human study, we trained the users to perform some actions depending on the chosen tactile feedback patterns. A human-robot collaborative task was conducted to evaluate the system, including the new design of our motion capture system (ArUcoGlide).
@@ 2 MoveTouch Motion Capture System
    A wearable robot, dubbed MoveTouch, that is responsible for adjusting the orientation of an ArUco marker held at its end-effector and providing haptic feedback to the user. The movable ArUco marker ensures the visibility of the marker to the motion capture system, avoiding any occlusion due to the movement of the user’s hand. The haptic feedback is provided to the user through five vibration motors controlled independently, which are able to generate different vibration patterns. The haptic system acts as a guidance system that tells the user about the optimal movement of their hand during Human-Robot Interaction (HRI) to avoid collisions.
@@ 2.1 MoveTouch Wearable Robot
The primary goal of the MoveTouch wearable robot is to assist in determining the user’s position and provide vibrotactile guidance, ensuring a secure HRI experience.
@@ 2.1.1 Mechanical Design
The proposed device is a two-degrees-of-freedom (2-DoFs) wearable robot, based on a differential gear train mechanism, that holds an ArUco marker at its end-effector with the ability to rotate this marker around two perpendicular axes through two servo motors [12], [14]. Additionally, it includes an array of vibration motors to deliver vibration patterns to the user’s wrist. The links, holders, and gears were designed and 3D-printed with PLA material; the 3D CAD model is shown in Fig. 1. The device maintains the visibility of the ArUco marker to the tracking system by continuously adjusting the motors’ angular position to hold the marker in a fixed orientation with respect to the camera coordinate system.
Figure 2: MoveTouch design: 3D model perspectives (a) isometric view, (b) top view revealing the gear structure.
To achieve the rotation of the end-effector around two perpendicular axes using identical motors positioned side by side, we decided to use a differential gear train mechanism, as shown in Fig. 1(a). In this figure, the fixed base, the differential gear train, and the rotating Aruco holder are shown.
We denote $\Delta\theta_{A}$ and $\Delta\theta_{B}$ as the rotation angles of the first and second motors, respectively. The rotation angles that the ArUco marker should undergo around the lateral and longitudinal axes are $\Delta\theta_{P}$ and $\Delta\theta_{C}$ which can be seen in Fig. 1(b). These angles, $\Delta\theta_{P}$ and $\Delta\theta_{C}$, can be acquired from the motion capture system to ensure the ArUco marker maintains a fixed orientation relative to the camera coordinate system [2]. The equations for determining the required rotation angles of both motors, based on the desired rotation angles for the marker, are as follows:
$\Delta\theta_{A}=n_{a}(n_{s}\Delta\theta_{P}+\Delta\theta_{C})$ (1)
@@
where $n_{a}=n_{b}=0.5$ are the gear ratio of the motor gears to the sun gears and $n_{s}=0.5$ is the gear ratio of the sun gears to the planet gear. The rotation angles $\Delta\theta_{A}$ and $\Delta\theta_{B}$ are acquired via the motion capture system.
We chose to integrate five vibration motors within the system. These motors were chosen due to their small size. The selection of this number takes into consideration two key factors: 1) Enabling a uniform distribution of the vibration motors on one side of the user’s wrist (either dorsal or volar), which typically falls within the range of 7.5 -10 cm. 2) This number of vibration motors is sufficient for generating a diverse array of vibrotactile patterns to be employed in user experiments.
@@ 2.1.2 Electronic Design
The electronic setup includes an ESP32 microcontroller along with two Gotech GS-9025MG servo motors. The system is powered by a Li-Po 7.4 V battery connected through a DC/DC converter to power the microcontroller, the servo driver and the vibration motors.
Five coin vibration motors, capable of vibrating at frequencies ranging from 10 to 55 Hz, are positioned on the bracelet with a 2 cm separation between them.
The angular position of the servo motors and the activation of the vibration motors are controlled via Bluetooth from a base computer.
@@ 2.2 Motion capture system
To ensure the safety of users, we need to track the real-time position of the operator within the working space. In this study, the interaction between one user wearing the MoveTouch robot on their wrist and a UR10 robotic manipulator was studied during collaborative task. Given that the MoveTouch marker is situated above the user’s hand as shown in Fig. 1, we suggest that determining the marker’s position enables us to locate the hand by applying a transformation from the marker to the center of the hand. In this paper, we will use the terms "user’s hand position" and "Movetouch marker" interchangeably.
To achieve a safe interaction, we need to locate the user’s hand in the robot coordinate system to make the robot avoid any collision with it. Our proposed mocap system is both cost-efficient and easy to install, utilizing an ArUco marker to track the operator’s hand. The system comprises a base computer and an RGB webcam C930e from Logitech mounted on a stand that can be adjusted to capture different angles of the workspace, providing greater flexibility for the user. The basic task of the motion capture system is to transfer the user’s hand position into the UR10 robot base coordinate system so that the robot can avoid collision with it. Since the MotionTouch marker, and thereby the user’s hand position, can be estimated in the camera coordinate system [8], the transformation from the MotionTouch marker to the camera $T_{C}^{A}$ can be known. By attaching another Aruco marker (we will call it the base marker) at a known transformation from the UR10 robot base, we can also estimate the position of the base of the robot in the camera coordinate system; thus, we can get the transformation $T_{C}^{B}$ from the UR10 robot base to the camera coordinate system. As a result, we obtain the transformation between the user’s hand and the UR10 robot base coordinate system $T_{B}^{A}$, and thus the position of the user’s hand in the UR10 robot base coordinate system as follows:
@@
Calculating the transformation from the base marker to the camera $T_{C}^{B}$ is required only once before starting the experiments, as the camera will be in the same position throughout the whole experiment. However, if the camera’s position or orientation is altered, the process needs to be repeated to derive the correct transformation matrix. Once the transformation matrix $T_{B}^{C}$ is determined, we can utilize it to track the position of the MotionTouch marker that is attached to the user’s hand, enabling us to locate it within the UR10 robot’s coordinate system.
@@ 3 Vibrotactile Guidance System
The proposed wearable device, MoveTouch, provides vibrotactile guidance, facilitating human-robot collaboration. This innovative system harnesses tactile sensations to guide users in adjusting their hand position when the collaborative robot approaches and alerts them about potentially risky situations. The system transfers information about how the user should move their hands through vibration patterns exerted on the user’s wrist. Ten vibration patterns have been chosen for testing in terms of recognition and discriminability on both the volar and dorsal parts of the user’s hand
@@ 3.1 Patterns Design
Considering that our wearable robot is worn on the wrist of the user, it is logical to incorporate vibrotactile patterns in this specific region. Our initial step is to evaluate the user’s ability to recognize vibrotactile patterns applied to their wrist and determine suitable frequencies for these patterns.
Five different patterns were designed using the five vibration motors, each with two different frequencies. The vibration patterns are illustrated in Fig. 3. In the first pattern, the vibration propagates sequentially through the five vibration patterns from the right side of the user to the left side. The second pattern is similar but the vibration propagation is from left to right. In the third pattern, vibration propagates symmetrically from the center to the outside. Initially, the central vibrator starts vibrating, followed by the simultaneous activation of the two neighboring vibrators. Subsequently, the vibration spreads to the two outermost vibrators. The same methodology for the fourth pattern but the direction is inversed to be from outside to the center. Finally, in the fifth pattern, all vibrators are activated together.
@@ 3.2 Pattern Recognition
In this experiment, we aimed to assess the ability of the participant to recognize and differentiate between the designed vibrotactile patterns. We invited 11 participants and asked them to tell us which haptic pattern they perceived. Their data was recorded and analyzed.
@@ 3.2.2 User Study Procedure
During the study, each user was asked to wear the MoveTouch robot and sit in front of a PC with a graphical user interface (GUI) that allowed them to choose the pattern that they felt. An example of the user study of this experiment is shown in Fig. 4(a).
Figure 4: The experimental setup: a) The pattern recognition experiment setup. b) The system evaluation in a real-world collaborative task setup.
To identify the most effective part form the wrist to perceive vibration (either dorsal or volar), participants were evaluated twice, wearing the device each time while the vibrators were either on the dorsal or volar part of their wrists. The participants were split into two groups: the first group tried the vibrators on the volar part first, followed by the dorsal part, while the second group followed the opposite sequence. Each of the 10 patterns was replicated five times in random order; thus, 50 patterns were provided to each participant in each evaluation. Participants were asked to wear headphones while playing white noise to prevent them from hearing the vibrations and engaged in the experiment through the GUI interface. They had the option to initiate the vibration pattern when they were ready and select from the displayed patterns the one they felt. During the experiment, participants were not allowed to repeat any pattern.
@@ 3.2.3 Results
From the two confusion matrices, the average recognition rate of all patterns on the dorsal part of the wrist was 66.6%, while on the volar side of the wrist, it was 75.64%.
@@ 3.2.4 Conclusion
We can notice that almost all the patterns are better recognized on the lower part of the wrist; thus, this position of the vibration motors was selected. These patterns include 1H, 1L, 2L, 3L, 4L, and 5H. However, to avoid user confusion, similar patterns like 1H and 1L are not included together. Instead, the pattern with the highest recognition, which is 1L, is selected. Consequently, patterns 1L, 2L, 3L, and 5H will used for further studies.
3.3 System Evaluation in a Collaborative Task
In this evaluation, the system was integrated into a real collaborative task to assess its effectiveness in enhancing the overall user safety and guiding users in a desired direction. The four selected patterns were integrated to direct users on hand movements as follows: Pattern 1L indicates moving the hand to the right, pattern 2L indicates moving the hand to the left, pattern 3L indicates moving the hand down, and pattern 5H indicates moving the hand back.
@@ 3.3.2 Experimental setup
The participants were asked to stand in front of a Siegmund welding table, where a UR10 robot was performing a pick-and-place task of cubes positioned at specific locations on the table. The participants wore the MoveTouch on their right wrist and utilized an eye cover to eliminate visual feedback about the location of the robot’s Tool Center Point (TCP). Participants were also performing a pick-and-place task in the same box that the robot was using. A surveillance camera was placed on a tripod near the table, where it could monitor the entire workspace. Additionally, An observer was present during the experiment, poised to activate the emergency button to halt the robot in case of any unforeseen issues. The experimental setup is shown in Fig. 4(b).